OpenStudy (nottim):

A reaction having time-independent stoichiometry follows the equation A + 2B → 2Z Suppose that initially there are present 0.28 moles if A, 0.39 moles of B and 0.13 moles of Z. After a certain amount of time, 0.18 moles of A remain. What amounts of B and Z are present.

OpenStudy (nottim):

Any ideas?

OpenStudy (michele_laino):

hint: at initial time, we have: \[\Large {n_A} = 0.28,\quad {n_B} = 0.39,\quad {n_C} = 0.13\] and at final time, the same number of moles become: \[\large {n_A} - x = 0.18,\quad {n_B} - x = 0.39 - 2x,\quad {n_C} + 2x = 0.13 + 2x\] so we can write: \[\Large 0.28 - x = 0.18\] Please solve for \(x\)